OpenStudy (anonymous):

Last onnee :DD x + 3y = -5 2x - y = 18 Solve the systems using substitution

OpenStudy (anonymous):

start with x=-5-3y and plug that into the 2nd equation.

OpenStudy (anonymous):

wanna try ?

OpenStudy (anonymous):

i would start with x=3(-18+2x)=-5

OpenStudy (anonymous):

that works too :)

OpenStudy (anonymous):

??? huh ill try but i do not understand :( and which one do i start with

OpenStudy (anonymous):

you can try doing it both ways for practice.

OpenStudy (anonymous):

you'll get the same answer either way

OpenStudy (anonymous):

using this method you need to find a way to plug in one of the equations into the other one you get that part right

OpenStudy (anonymous):

x=3(-18+2x)=-5 x= -54+6x=-5 +54 +54 x=6x-5 ???? i dont think im doing this right haha

OpenStudy (anonymous):